# Maximizing the Azimuthal-Angle Correlation in the Decay of Vector Boson Pairs

### v1

@@ Abstract
Spin correlation observables of particles produced at colliders generally depend on the coordinate basis. While spin correlations and their associated basis dependence have been studied in top-quark pair production, the basis dependence of spin correlations of a $W^{\pm}$ pair has not been studied. In this work, we focus on the azimuthal-angle correlation in the decay of a $W^{\pm}$ pair, which encodes spin information in the direction transverse to the rotational axis. We quantitatively describe the basis dependence of the relative azimuthal angle. We further present the optimal coordinate direction to define the azimuthal angle between the decay products of $W^{\pm}$ bosons to achieve the maximal spin correlation based on symmetry considerations. The results can be used to optimize the measurement of the CP-even and CP-odd interactions, and have applications to quantum information observables in the vector boson pair system.
@@ II Angular Correlations in Decays and its Basis Dependence
$\hat{k}=\hat{p}_{W^{-}},\qquad\qquad\hat{r}=\frac{\hat{p}_{e^{-}}-\hat{k}\cos\Theta}{\sin\Theta},\qquad\qquad\hat{n}=\hat{k}\times\hat{r},$ (12)
where $\hat{p}_{e^{-}}/\hat{p}_{W^{-}}$ is the normalized three-momentum direction of $e^{-}/W^{-}$, and $\Theta$ is the scattering angle between $e^{-}$ and $W^{-}$, as shown in Fig. 3(a). The $\Delta\phi$ distributions with respect to these three rotational axes are shown in Fig. 3(b), and we see that the distribution exhibits a strong dependence on the choice of the rotational axis. This offers an opportunity to identify a reference rotational axis to optimize the azimuthal-angle correlation. We also find that the higher-order terms $\cos(2\Delta\phi)$ and $\sin(2\Delta\phi)$ are negligible, and we thus only focus on the two leading terms hereafter.
@@ IV Conclusions
Finally, we note that there is potential for future work when the $W^{+}W^{-}$ system is treated as a quantum state. While our work optimizes a single entry of the spin correlation matrix, analogous to Refs. [1, 2, 3] in the $t\bar{t}$ system, a natural follow-up would be to perform optimization for quantum entanglement or Bell nonlocality, as was done in Refs. [4, 5] for the $t\bar{t}$ system.
@@ Appendix A The Production Density Matrix and Spin Correlation Matrix of a $W^{\pm}$ Boson Pair
$\mathcal{N}=\sum_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\left|\mathcal{M}^{\rm SM}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}+\mathcal{M}^{\rm EFT}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\right|^{2}=\sum_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\left|\mathcal{M}^{\rm EFT}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\right|^{2}+\sum_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\left|\mathcal{M}^{\rm EFT}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\right|^{2}.$ (35)

### v2

@@ Abstract
Spin correlation observables of particles produced at colliders generally depend on the coordinate basis. While spin correlations and their associated basis dependence have been studied in top-quark pair production, the basis dependence of spin correlations of $W^{\pm}$ pairs has not been studied. In this work, we focus on the azimuthal-angle correlation in the decay of $W^{\pm}$ pairs, which encodes spin information in the direction transverse to the rotational axis. We quantitatively describe the basis dependence of the relative azimuthal angle. We further present the optimal coordinate direction to define the azimuthal angle between the decay products of $W^{\pm}$ bosons to achieve the maximal spin correlation based on symmetry considerations. The results can be used to optimize the measurement of the CP-even and CP-odd interactions, and have applications to quantum information observables in the vector boson pair system.
@@ II Angular Correlations in Decays and its Basis Dependence
$\hat{k}=\hat{p}_{W^{-}},\qquad\qquad\hat{r}=\frac{\hat{k}\cos\Theta-\hat{p}_{e^{-}}}{\sin\Theta},\qquad\qquad\hat{n}=\hat{k}\times\hat{r},$ (12)
where $\hat{p}_{e^{-}}/\hat{p}_{W^{-}}$ is the normalized three-momentum direction of $e^{-}/W^{-}$, and $\Theta$ is the scattering angle between $e^{-}$ and $W^{-}$, as shown in Fig. 3(a). The $\Delta\phi$ distributions with respect to these three rotational axes are shown in Fig. 3(b), and we see that the distribution exhibits a strong dependence on the choice of the rotational axis. This offers an opportunity to identify a reference rotational axis to optimize the azimuthal-angle correlation. The dashed curves in Fig. 3(b) only include the leading term in the $\Delta\phi$ distribution. We find that the higher-order terms, $\cos(2\Delta\phi)$ and $\sin(2\Delta\phi)$, are negligible and we only focus on the leading terms hereafter.
@@ III.1 Optimizing the Azimuthal-Angle Correlation: CP-Even Interactions
The above optimal basis is obtained at leading order. A complete higher-order treatment is beyond the scope of this work. However, since the NLO electroweak correction to the spin observables of $e^{+}e^{-}\to W^{+}W^{-}$ is within 10% [23], we estimate that the optimal basis at tree level still serves as the approximately optimal basis at higher orders, especially if any additional high $p_{T}$ objects in the events are vetoed to explicitly maintain the leading order kinematics.
@@ IV Conclusions
Finally, we note that there is potential for future work when the $W^{+}W^{-}$ system is treated as a quantum state (see also Ref. [27]). While our work optimizes a single entry of the spin correlation matrix, analogous to Refs. [1, 2, 3] in the $t\bar{t}$ system, a natural follow-up would be to perform optimization for quantum entanglement or Bell nonlocality, as was done in Refs. [4, 5] for the $t\bar{t}$ system.
@@ Appendix A The Production Density Matrix and Spin Correlation Matrix of a $W^{\pm}$ Boson Pair
$\mathcal{N}=\sum_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\left|\mathcal{M}^{\rm SM}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}+\mathcal{M}^{\rm EFT}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\right|^{2}=\sum_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\left|\mathcal{M}^{\rm SM}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\right|^{2}+\sum_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\left|\mathcal{M}^{\rm EFT}_{\sigma\bar{\sigma},\lambda\bar{\lambda}}\right|^{2}.$ (35)